the sum of two integers is 42. their difference is 8 find the integers.
Answer by Theo(13342)

let x = one of the integers
let y = the other of the integers

The sum is 42 gets you x + y = 42

The difference is 8 gets you x - y = 8

You need to solve these 2 equations simultaneously to get your answers.

The equations are:

x + y = 42
x - y = 8

Add the 2 equations together to get:

2x = 50

Divide both sides of this equation by 2 to get:

x = 25


If x = 25, then x + y = 42 gets you y = 17

If x = 25 and y = 17, then x - y = 8 gets you 25 - 17 = 8 which becomes 8 = 8 which is true.

This confirms that the values of 25 for x and 17 for y are good.

Your answer is that the integers are 25 and 17.

25 + 17 = 42
25 - 17 = 8